# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 3 Sequences of bifurcations: geometric method of determination
The Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ in Fig. 1 were obtained by fixing the value of the energy $\mathcal{E}$ and taking a set of initial conditions for each energy. Hence, every periodic orbit $P$ yields a fixed point with coordinates $X_{2,P}(\mathcal{E}),Y_{2,P}(\mathcal{E})$, with $\mathcal{E}$, in general, within two limits $\mathcal{E}_{min,P}\leq\mathcal{E}\leq\mathcal{E}_{max,P}$. The energies $\mathcal{E}_{min,P},\mathcal{E}_{max,P}$ are called the bifurcation limits of the periodic orbit $P$, while the curves $C_{P}=\mbox{graph of\leavevmode\nobreak\ }X_{2,P}(\mathcal{E}),Y_{2,P}(\mathcal{E})$ as a function of $\mathcal{E}$ are called characteristic curves.
@@
Consider all possible periodic orbits $P_{n}$, $n=1,2,\ldots$ yielding fixed points on the section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ of a given character of Floquet stability $S_{n}$ (with values ‘stable’ or ’unstable’), and sorted in ascending values of $\mathcal{E}_{min,P_{n}}$. The set $\mathcal{B}\equiv\{(\mathcal{E}_{min,P_{n}},\mathcal{E}_{max,P_{n}},C_{P_{n}},S_{n}),n=1,2,\ldots\}$ is hereafter called a bifurcation sequence.
@@
The definition of a bifurcation sequence $\mathcal{B}$ as above requires parameterizing all bifurcation limits in terms of the corresponding energies $(\mathcal{E}_{min,P_{n}},\mathcal{E}_{max,P_{n}})$. To predict the structural changes in the corresponding phase portraits after the n-th bifurcation in the sequence, we employ integrable approximations to the dynamics, as the models $\mathcal{H}_{int}$, $\mathcal{H}_{int}^{(1)}$, or ~$\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$ mentioned in the introduction. An integrable approximation $\mathcal{Z}$ to the Hamiltonian $\mathcal{H}_{sec}$ is defined as a Hamiltonian which i) admits $\sigma_{0}$ (Eq.(4)) as a second integral, and ii) has flow close to the flow of $\mathcal{H}_{sec}$. The closeness of the two flows can be established by appropriate norms in the functional space containing the derivatives of $\mathcal{H}_{sec}$ and $\mathcal{Z}$. Let $\mathcal{Z}$ be an integrable approximation to $\mathcal{H}_{sec}$. Define the family of curves
[…]
$X_{2}^{2}+Y_{2}^{2}+X_{3,+}^{2}=2\sigma_{0},\leavevmode\nobreak\ \dot{Y}_{3}=-{\partial\mathcal{Z}\over\partial X_{3}}\big{|}_{Y_{3}=0,X_{3}=X_{3,+}}\geq 0,$ (19)
$\cos(\mathrm{i}_{max})\leq\cos(\mathrm{i}_{mut})(X_{2},Y_{2},X_{3}=X_{3,+},Y_{3}=0;\mathrm{AMD})\leq 1\bigg{\}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$
@@
A $\sigma_{0}-$fixed phase portrait of $\mathcal{Z}$ is defined by the family of curves
$\mathcal{P_{Z}}(\sigma_{0})=\bigg{\{}\mathcal{C_{Z}}(\sigma_{0};\mathcal{E};\mathrm{AMD}),\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{E}_{L}(\sigma_{0},\mathrm{AMD})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0},\mathrm{AMD})\bigg{\}}\,,$ (20)
@@ A $\sigma_{0}-$fixed phase portrait of $\mathcal{Z}$ is defined by the family of curves
$\mathcal{P_{Z}}(\mathcal{E})=\left\{\mathcal{C_{Z}}(\sigma_{0};\mathcal{E};\mathrm{AMD}),\leavevmode\nobreak\ \leavevmode\nobreak\ \sigma_{0,min}(\mathcal{E},\mathrm{AMD})\leq\sigma_{0}\leq\sigma_{0,max}(\mathcal{E},\mathrm{AMD})\right\}$ (21)
@@
Figure 3: The permissible domain in $\mathcal{E}$ and $\sigma_{0}$ consistent with a fixed value of the system’s $\mathrm{AMD}$. The numerical curves shown above refer to the example of the integrable model $\mathcal{Z}=\mathcal{H}_{int}$. Fixing a value of $\sigma_{0}$, the two limiting curves yield the corresponding limiting energies $\mathcal{E}_{L}(\sigma_{0})$, $\mathcal{E}_{R}(\sigma_{0})$. Alternatively, fixing a value of the energy we obtain the limits $\sigma_{0,min}(\mathcal{E}),\sigma_{0,max}(\mathcal{E})$. The above limits can be computed analytically as explained in [13]. In summary, there exist two critical values of the energy $\mathcal{E}_{min}$, $\mathcal{E}_{2,3}$ such that, for any value of the energy in the interval $\mathcal{E}_{min}<\mathcal{E}<\mathcal{E}_{2,3}$, $\sigma_{0}$ is limited from below by a minimum value $\sigma_{0,min}(\mathcal{E})$, while the limit from above $\sigma_{0,max}$ is posed only by the globally maximum value $\sigma_{0,max}=\sigma_{0,\mathrm{AMD}}=\mathrm{AMD}$. The limiting values $\mathcal{E}=\mathcal{E}_{min}$ and $\mathcal{E}=\mathcal{E}_{2,3}$ correspond to the co-planar ACR orbits A (anti-aligned) and B (aligned) respectively. On the other hand, for energies larger than $\mathcal{E}_{2,3}$ there are no possible co-planar orbits, and $\sigma_{0,max}$ as well becomes a decreasing function of $\mathcal{E}$. Overall, as we move from left to right or top to bottom, we obtain orbits of higher mutual inclination. For the Hamiltonian $\mathcal{H}_{int}$ (see text for parameter values) we compute the values $\mathcal{E}_{min}=-1.18\cdot 10^{-4}$ and $\mathcal{E}_{2,3}=-6.77\cdot 10^{-5}$ and $\sigma_{0_{max}}=\mathrm{AMD}=0.0162044$.
@@
i) $\mathcal{E}-$fixed phase portraits of $\mathcal{Z}$ are equivalent to the Poincaré surfaces of section $\mathcal{PS}_{\mathcal{Z}}(\mathcal{E};\mathrm{AMD})$. The equivalence is demonstrated below. Since the flow of $\mathcal{Z}$ is close to the flow of $\mathcal{H}_{sec}$ in a suitable domain of initial conditions the phase portraits $\mathcal{P}_{\mathcal{Z}}(\mathcal{E})$ are expected to approximate as well the phase portraits of the full non-integrable Hamiltonian $\mathcal{H}_{sec}$, obtained through the Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$.
ii) For fixed integrable model $\mathcal{Z}$, the phase portraits $\mathcal{P_{Z}}(\mathcal{E})$ and $\mathcal{P_{Z}}(\sigma_{0})$ yield an equivalent sequence of bifurcations. This is an obvious consequence of the definition of bifurcation sequence together with the fact that the sets $\mathcal{P_{Z}}(\mathcal{E})$ and $\mathcal{P_{Z}}(\sigma_{0}$, spanning the whole permissible domain in $\mathcal{E}$ and $\sigma_{0}$ (Fig.3), contain the same curves $\mathcal{C_{Z}}(\sigma_{0},\mathcal{E};\mathrm{AMD})$. An example of this equivalence is provided in Fig. 2.
Note that, due to the last property mentioned in (i), only the representation $\mathcal{P}_{\mathcal{Z}}(\mathcal{E})$ can be compared directly with the numerical phase-portraits under the full (non-integrable) Hamiltonian $\mathcal{H}_{sec}$. On the other hand, due to (ii), the theoretical analysis of the bifurcation sequences can be performed exploring the geometric properties of the representation $\mathcal{P_{Z}}(\sigma_{0})$ in suitable Hopf variables (see next subsection). The equivalence property then establishes the equality of the bifurcation sequences exhibited by either representation.
@@
$\psi=w_{2}-w_{3}\,,$ $\Gamma=\frac{W_{2}-W_{3}}{2}\,,$
$\varphi=w_{2}+w_{3}\,,$ $J=\frac{W_{2}+W_{3}}{2}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (22)
@@
$\mathcal{Z}(\psi,\varphi,\Gamma,J)=\mathcal{Z}(\psi,\Gamma;J)\leavevmode\nobreak\ \leavevmode\nobreak\ ,$ (23)
@@
i.e., the angle $\varphi$ is ignorable. In view of Eq. (23), $\mathcal{Z}$ can be regarded as defining either the complete flow in four variables $(\psi,\varphi,\Gamma,J)$, or, alternatively, the flow a one degree of freedom Hamiltonian in the variables $(\psi,\Gamma)$, with $J$ serving as parameter. Hence, the level curves of constant energy $\mathcal{Z}(\psi,\Gamma;J)=\mathcal{E}$ computed in the plane $(\psi,\Gamma)$ for various values of the energy and keeping fixed the value of $J$ are geometrically equivalent to the invariant curves in the Poincaré section under the complete 2-DOF flow. In particular, any closed curve $\gamma$ (1-torus) in the top frame of Fig. 2 represents a 2-torus of the full flow, obtained by the product $\gamma\times$ the 1-torus defined by the solution for $\varphi$ of the differential equation $\dot{\varphi}=(\partial\mathcal{Z}/\partial J)_{\psi_{\gamma}(t),\Gamma_{\gamma}(t)}$, where $\psi_{\gamma}(t),\Gamma_{\gamma}(t)$ is the solution of the Hamiltonian flow for the variables $(\psi,\Gamma)$ along the curve $\gamma$. In the same way, any fixed point $(\psi_{0},\Gamma_{0})$ in the same plots represents a periodic orbit $\varphi(t)=\varphi(0)+(\partial\mathcal{Z}/\partial J)_{\psi=\psi_{0},\Gamma=\Gamma_{0}}$ of the complete flow. Hence, the problem of the number and bifurcations of new periodic orbits in the complete 2-DOF flow is reduced to the problem of the number and bifurcations of new fixed points in the reduced 1-DOF flow.

### v3

@@ 3 Sequences of bifurcations: geometric method of determination
The Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ in Fig. 1 were obtained by fixing the value of the energy $\mathcal{E}$ and taking a set of initial conditions for each energy.
Definition 3.1.
@@
Let $P$ be a periodic orbit of the flow under $\mathcal{H}_{sec}$ at the energy value $\mathcal{H}_{sec}=\mathcal{E}$. To the orbit $P$ corresponds a fixed point in the Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ with coordinates $X_{2,P}(\mathcal{E}),Y_{2,P}(\mathcal{E})$. The orbit $P$ exists, in general, within two energy limits $\mathcal{E}_{min,P}\leq\mathcal{E}\leq\mathcal{E}_{max,P}$. We call $\mathcal{E}_{min,P}$ and $\mathcal{E}_{max,P}$ the bifurcation limits of the periodic orbit $P$. The graphs of $X_{2,P}(\mathcal{E})$, $Y_{2,P}(\mathcal{E})$ as a function of $\mathcal{E}$ are called characteristic curves $C_{P}$ of the periodic orbit $P$.
The procedure to compute the bifurcation limits of particular periodic orbits is detailed in Section $2.4$ of [17]. In particular, a fixed value of $\mathrm{AMD}$ determines the maximal allowed value of the mutual inclination for any planetary orbits of the system. Consider the domain in mutual inclination $0,\mathrm{i}_{mut_{max}}]$. A certain value of the energy $\mathcal{E}$ establishes, within that domain, a range of allowed mutual inclinations (see figure 8 of [17]). To find it, it is sufficient to look for tangencies between the manifold of constant energy to the section $Y_{3}=0$, i.e.
$\mathcal{M}(\mathcal{E})=\bigg\{(X_{2},Y_{2},X_{3})\in\mathbb{R}^{3}:\mathcal{H}_{sec}(X_{2},X_{3},Y_{2},Y_{3}=0;\mathrm{AMD})=\mathcal{E}\bigg\}$
and the ellipsoidal surface
$\mathcal{I}_{\mathrm{i}_{mut}}=\bigg\{(X_{2},Y_{2},X_{3})\in\mathbb{R}^{3}:\frac{L_{z}^{2}-\Lambda_{2}^{2}\left(1-{X_{2}^{2}+Y_{2}^{2}\over 2\Lambda_{2}}\right)^{2}-\Lambda_{3}^{2}\left(1-{X_{3}^{2}\over 2\Lambda_{3}}\right)^{2}}{2\Lambda_{2}\Lambda_{3}\left(1-{X_{2}^{2}+Y_{2}^{2}\over 2\Lambda_{2}}\right)\left(1-{X_{3}^{2}\over 2\Lambda_{3}}\right)}=\cos(\mathrm{i}_{mut})\bigg\}\penalty 10000\ \penalty 10000\ ,$
with angular momentum $L_{z}=\Lambda_{2}+\Lambda_{3}-\mathrm{AMD}$ (see Eq. (29) of [17]). Once established, the correspondence between energies $\mathcal{E}$ and mutual inclinations $\mathrm{i}_{mut}$ allows to find the bifurcation limits $\mathcal{E}_{min,P}$, $\mathcal{E}_{max,P}$ for a periodic orbit $P$.
Figure 2: The permissible domain in $\mathcal{E}$ and $\sigma_{0}$ consistent with a fixed value of the system’s $\mathrm{AMD}$. The numerical curves shown above refer to the example of the integrable model $\mathcal{Z}=\mathcal{H}_{int}$. Fixing a value of $\sigma_{0}$, the two limiting curves yield the corresponding limiting energies $\mathcal{E}_{L}(\sigma_{0})$, $\mathcal{E}_{R}(\sigma_{0})$. Alternatively, fixing a value of the energy we obtain the limits $\sigma_{0,min}(\mathcal{E}),\sigma_{0,max}(\mathcal{E})$. The above limits can be computed analytically as explained in [17] (computing the tangencies between the sphere of radius $\sigma_{0}$ described in (33) and the constant energy surface reported in (34)). In summary, there exist two critical values of the energy $\mathcal{E}_{min}$, $\mathcal{E}_{2,3}$ such that, for any value of the energy in the interval $\mathcal{E}_{min}<\mathcal{E}<\mathcal{E}_{2,3}$, $\sigma_{0}$ is limited from below by a minimum value $\sigma_{0,min}(\mathcal{E})$, while the limit from above $\sigma_{0,max}$ is posed only by the globally maximum value $\sigma_{0,max}=\sigma_{0,\mathrm{AMD}}=\mathrm{AMD}$. The limiting values $\mathcal{E}=\mathcal{E}_{min}$ and $\mathcal{E}=\mathcal{E}_{2,3}$ correspond to the co-planar ACR orbits A (anti-aligned) and B (aligned) respectively. On the other hand, for energies larger than $\mathcal{E}_{2,3}$ there are no possible co-planar orbits, and $\sigma_{0,max}$ as well becomes a decreasing function of $\mathcal{E}$. Overall, as we move from left to right or top to bottom, we obtain orbits of higher mutual inclination. For the Hamiltonian $\mathcal{H}_{int}$ (see text for parameter values) we compute the values $\mathcal{E}_{min}=-1.18\cdot 10^{-4}$ and $\mathcal{E}_{2,3}=-6.77\cdot 10^{-5}$ and $\sigma_{0_{max}}=\mathrm{AMD}=0.0162044$.
Definition 3.2.
@@
Consider all possible periodic orbits $P_{n}$, $n=1,2,\ldots$, yielding fixed points on the section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$ of a given character of Floquet $S_{n}$ (with values ‘stable’ or ’unstable’), and sorted in ascending values of $\mathcal{E}_{min,P_{n}}$. The set $\mathcal{B}\equiv\{(\mathcal{E}_{min,P_{n}},\mathcal{E}_{max,P_{n}},C_{P_{n}},S_{n}),n=1,2,\ldots\}$ is hereafter called a bifurcation sequence.
The periodic orbit $P_{n}$ is called Floquet unstable if the Floquet matrix of the linearized flow around the periodic orbit has at least one Floquet exponent with positive real part. Otherwise, the periodic orbit is called Floquet stable (see [24]).
@@
The definition of a bifurcation sequence $\mathcal{B}$ as above requires parameterizing all bifurcation limits in terms of the corresponding energies $(\mathcal{E}_{min,P_{n}},\mathcal{E}_{max,P_{n}})$. To predict the structural changes in the corresponding phase portraits after the n-th bifurcation in the sequence, we employ integrable approximations to the dynamics, as the models $\mathcal{H}_{int}$, $\mathcal{H}_{int}^{(1)}$, or $\tilde{\mathcal{H}}_{int}^{(N_{P}=3,N_{bk}=4)}$ mentioned in the introduction. An integrable approximation $\mathcal{Z}$ to the Hamiltonian $\mathcal{H}_{sec}$ is defined as a Hamiltonian which i) admits $\sigma_{0}$ (Eq.(4)) as a second integral, and ii) has flow close to the flow of $\mathcal{H}_{sec}$. The closeness of the two flows can be established by appropriate norms in the functional space containing the derivatives of $\mathcal{H}_{sec}$ and $\mathcal{Z}$.
Definition 3.3.
Let $\mathcal{Z}$ be an integrable approximation to $\mathcal{H}_{sec}$. Define the family of curves
$\mathcal{C_{Z}}(\sigma_{0};\mathcal{E};\mathrm{AMD})=\bigg\{$ $(X_{2},Y_{2})\in\mathbb{R}^{2}:\penalty 10000\ \mathcal{Z}(X_{2},Y_{2},X_{3}=X_{3,+},Y_{3}=0;\mathrm{AMD})=\mathcal{E},$
$X_{2}^{2}+Y_{2}^{2}+X_{3,+}^{2}=2\sigma_{0},\penalty 10000\ \dot{Y}_{3}=-{\partial\mathcal{Z}\over\partial X_{3}}\big|_{Y_{3}=0,X_{3}=X_{3,+}}\geq 0,$ (23)
$\cos(\mathrm{i}_{max})\leq\cos(\mathrm{i}_{mut})(X_{2},Y_{2},X_{3}=X_{3,+},Y_{3}=0;\mathrm{AMD})\leq 1\bigg\}\penalty 10000\ \penalty 10000\ ,$
@@
where $X_{3,+}=\sqrt{2\sigma_{0}-X_{2}^{2}-Y_{2}^{2}}\,$. A $\sigma_{0}$-fixed phase portrait of $\mathcal{Z}$ is defined by the family of curves
$\mathcal{P_{Z}}(\sigma_{0})=\bigg\{\mathcal{C_{Z}}(\sigma_{0};\mathcal{E};\mathrm{AMD}),\penalty 10000\ \penalty 10000\ \mathcal{E}_{L}(\sigma_{0},\mathrm{AMD})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0},\mathrm{AMD})\bigg\}\,$, (24)
@@
$\mathcal{P_{Z}}(\mathcal{E})=\left\{\mathcal{C_{Z}}(\sigma_{0};\mathcal{E};\mathrm{AMD}),\penalty 10000\ \penalty 10000\ \sigma_{0,min}(\mathcal{E},\mathrm{AMD})\leq\sigma_{0}\leq\sigma_{0,max}(\mathcal{E},\mathrm{AMD})\right\}$ (25)
@@
Proposition 3.4.
$\mathcal{E}$-fixed phase portraits of $\mathcal{Z}$, i.e. $\mathcal{P_{Z}}(\mathcal{E})$, are equivalent to the Poincaré surfaces of section $\mathcal{PS}_{\mathcal{Z}}(\mathcal{E};\mathrm{AMD})$.
Proof.
@@
$\psi=w_{2}-w_{3}\,$, $\Gamma=\frac{W_{2}-W_{3}}{2}\,$, (26)
$\varphi=w_{2}+w_{3}\,$, $J=\frac{W_{2}+W_{3}}{2}\penalty 10000\ \penalty 10000\ .$
@@
$\mathcal{Z}(\psi,\varphi,\Gamma,J)=\mathcal{Z}(\psi,\Gamma;J)\penalty 10000\ \penalty 10000\ ,$ (27)
@@
i.e., the angle $\varphi$ is ignorable. In view of Eq. (27), $\mathcal{Z}$ can be regarded as defining either the complete flow in four variables $(\psi,\varphi,\Gamma,J)$, or, alternatively, the flow of a one degree of freedom Hamiltonian in the variables $(\psi,\Gamma)$, with $J$ serving as parameter. Hence, the level curves of constant energy $\mathcal{Z}(\psi,\Gamma;J)=\mathcal{E}$ computed in the plane $(\psi,\Gamma)$ for various values of the energy and keeping fixed the value of $J$ are geometrically equivalent to the invariant curves in the Poincaré section under the complete 2-DOF flow. In particular, any closed curve $\gamma$ (1-torus) in the top frame of Fig. 3 represents a 2-torus of the full flow, obtained by the product $\gamma\times$ the 1-torus defined by the solution for $\varphi$ of the differential equation $\dot{\varphi}=(\partial\mathcal{Z}/\partial J)_{\psi_{\gamma}(t),\Gamma_{\gamma}(t)}$, where $\psi_{\gamma}(t),\Gamma_{\gamma}(t)$ is the solution of the Hamiltonian flow for the variables $(\psi,\Gamma)$ along the curve $\gamma$. In the same way, any fixed point $(\psi_{0},\Gamma_{0})$ in the same plots represents a periodic orbit $\varphi(t)=\varphi(0)+(\partial\mathcal{Z}/\partial J)_{\psi=\psi_{0},\Gamma=\Gamma_{0}}$ of the complete flow. Hence, the problem of the number and bifurcations of new periodic orbits in the complete 2-DOF flow is reduced to the problem of the number and bifurcations of new fixed points in the reduced 1-DOF flow. ∎
Note that, in the above proposition, ‘equivalence’ means that each member of the family of the 2D tori of the Hamiltonian flow of $\mathcal{Z}$, regarded as a function of four variables, can be mapped one-to-one to a member of the family of contours $\gamma(\mathcal{E},J)$.
Observation 3.1.
Since the flow of $\mathcal{Z}$ is close to the flow of $\mathcal{H}_{sec}$ in a suitable domain of initial conditions the phase portraits $\mathcal{P}_{\mathcal{Z}}(\mathcal{E})$ are expected to approximate as well the phase portraits of the full non-integrable Hamiltonian $\mathcal{H}_{sec}$, obtained through the Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{sec}}(\mathcal{E};\mathrm{AMD})$.
Observation 3.2.
For fixed integrable model $\mathcal{Z}$, the phase portraits $\mathcal{P_{Z}}(\mathcal{E})$ and $\mathcal{P_{Z}}(\sigma_{0})$ yield an identical sequence of bifurcations.
This is an obvious consequence of the definition of bifurcation sequence together with the fact that the sets $\mathcal{P_{Z}}(\mathcal{E})$ and $\mathcal{P_{Z}}(\sigma_{0}{)}$, spanning the whole permissible domain in $\mathcal{E}$ and $\sigma_{0}$ (Fig. 2), contain the same curves $\mathcal{C_{Z}}(\sigma_{0},\mathcal{E};\mathrm{AMD})$. An example is provided in Fig. 3.
Refer to captionRefer to caption
Figure 3: Top frame: Poincaré surfaces of section $\mathcal{PS}_{\mathcal{H}_{int}}(\mathcal{E};\mathrm{AMD})$ for the integrable Hamiltonian model $\mathcal{H}_{int}$ at the values of the energies (from top to bottom, left to right) $\mathcal{E}=-6.77\cdot 10^{-5},-1.8\cdot 10^{-5},-1.74\cdot 10^{-5},-1.7\cdot 10^{-5},-1.67\cdot 10^{-5},-1.6\cdot 10^{-5}$$1.43\cdot 10^{-5},-4.05\cdot 10^{-7}$. Bottom frame: Phase portraits by contour plots $\mathcal{P}_{\mathcal{H}_{int}}(\sigma_{0};\mathrm{AMD})$ with decreasing values of $\sigma_{0}$ (from top to bottom, left to right) $\sigma_{0}=1.62\cdot 10^{-2},6.15\cdot 10^{-3},5.93\cdot 10^{-3},5.85\cdot 10^{-3},5.78\cdot 10^{-3},5.62\cdot 10^{-3},5.304\cdot 10^{-3},1.74\cdot 10^{-4}$. The two alternative representations of the phase portraits, corresponding to $\mathcal{P_{Z}}(\mathcal{E})$ (top) or $\mathcal{P_{Z}}(\sigma_{0})$ (bottom), for $\mathcal{Z}=\mathcal{H}_{int}$, yield an equivalent bifurcation sequence.
Note that, due to Observation 3.1, only the representation $\mathcal{P}_{\mathcal{Z}}(\mathcal{E})$ can be compared directly with the numerical phase-portraits under the full (non-integrable) Hamiltonian $\mathcal{H}_{sec}$. On the other hand, due to Observation 3.2, the theoretical analysis of the bifurcation sequences can be performed exploring the geometric properties of the representation $\mathcal{P_{Z}}(\sigma_{0})$ in suitable Hopf variables (see next subsection). Observation 3.2 then implies that the two different representations of the phase portraits show an identical sequence of bifurcations, namely $\mathcal{B}$.